Determine all the roots of the given function below using Bairstow's Method with r=s=0 as initial values and Er=Es<=0.88154% as terminating conditions

f(z) = 2.3-12.2 + 32,-1
Answer #1

ク r-2 +2 Solve bo → 오 Tool s. otal 3 aoofs q! 叶(0) = 3 스叹 0,3846 AS-5/132 3 26683 o 3846 O.11632 2213030.bo 6 ss=-o-6970 0.85 S SHAS = 0.3846+(-6,990 660 Solut om o rt ve.gsJY-牛(6.6os4) χ=